Due Date Points Possible: Grade Category: Description: ing Learning 100 Astronomers discover an exoplanet (a planet of a star other than the Sun) that has an orbital period of 3.97 Earth years in its circular orbit around its sun, which is a star with a measured mass of 3.41 x1020 kg. Find the radius of the exoplanets orbit. Graded Policies: good You can check your answers You can view solutions when you complete ar give up on any question You have ten attempts per question There is no penalty for incarrect answers O Help With This Topic ○Web Help & O Technical Support and Bug Next Exit
Answer #1

Given,

orbital period of exoplanet = 3.97 Earth years

mass, m = 3.41 × 1030 kg

The orbital period of the planet = 3.97 earth years = 3.97 x 365 x 24 x 3600 seconds = 125197920 seconds

G = 6.6726 x 10-11N-m2/kg2

From kepler's third law equation

The orbital radius, r = (T2 GM/4π2)(1/3) = [(1251979202 x 6.67 x 10-11 x 3.41 x 1030 )/(4π2)]1/3 = 4.486 x 1011 m

Please rate my answer if you find it helpful, good luck...